# Taking quantisation seriously: a farewell to waves

### v2

@@ Abstract
The dual wave-particle nature of quantum objects is a notoriously unintuitive feature of quantum theories. However, it is often deemed essential, due to otherwise discrete particles exhibiting diffraction and interference. Troublingly, Landé and Lévy-Leblond have shown that de Broglie wavelengths are not relativistically covariant as simultaneous spatial structures, making any such wave properties physically inconsistent. In this work, we explore whether modern experiments vindicate an alternative view: that apparent waviness in diffraction and interference scenarios emerges as a consequence of quantised interactions between particles. Such a view has historically received very little attention, despite being the exact modern explanation of both the Kapitza-Dirac effect and ultrafast electron diffraction. We then study a photon orbital angular momentum realisation of the double slit to prove this explanation capable of unifying barrier and phase-grating diffraction phenomena. Finally, we demonstrate that the quantum formalism demands that particle momentum is determined at the point of scattering, contravening wave accounts of quantum interference.
@@ 1 Introduction
There is, however, a far graver problem with quantum waviness: the wavelengths are not covariant as correlation lengths in any form of relativity. In 1975 Landé argued [6] that because de Broglie wavelengths take the form $\lambda=\frac{h}{p}$ they must vary with inertial frame. Landé also assumed that all real waves should possess phase invariance, generating an apparent paradox. Lévy-leblond [7] showed that $\lambda$ varying with inertial frame is necessary to maintain the Galilean invariance of Schrödinger equation solutions and thus phase invariance does not hold for quantum waves. This establishes that $\lambda=\frac{h}{p}$ is not a Galilean invariant, indeed it is manifestly not covariant as a length in special relativity either. By this we mean that we cannot measure $\lambda$ as the simultaneous length of an extended structure in space. This is not the case for sound or water waves. So $\lambda$ cannot be interpreted as a simultaneous physical length scale. The solution in relativistic quantum theories is to rely on $p^{\mu}=\hbar k^{\mu}$ as a definition of a covariant wave-vector. This appears to leave us with a paradox. We can provide a covariant mathematical description, but the wavelength is not covariant in the manner we should expect. We should clearly ask what kind of covariance does $\lambda$ then exhibit. In answer, we consider the Lorentz transform for distances propagated by an object with speed $u\to c$, in some reference frame $S$, over a time $\Delta t$ or
$\Delta x^{\prime}=\Delta x\sqrt{\frac{1-\beta}{1+\beta}}\;,$ (1)
where $\beta$ is the speed of frame $S^{\prime}$ relative to $S$. The transformation has the same structure as the relativistic Doppler shift for electromagnetic wavelengths, providing an indication that such wavelengths transform like photon propagation distances, not the correlation length of waves extended in space. Moreover, a Doppler-like factor occurs similarly for any propagation speed $u$: $\Delta x^{\prime}=\gamma\Delta x\frac{(u-v)}{u}$. Further, we can express $p^{\prime}=\gamma p\frac{u-v}{u}$ for particle moving with $u\leq c$, which yields $\lambda^{\prime}=\frac{\lambda}{\gamma}\frac{u}{u-v}$, with the Galilean form being $\lambda^{\prime}=\frac{h}{m(u-v)}=\lambda\frac{u}{u-v}$. Repeating our pattern of similar structure to the transformation of propagation distances. Notably, our wave-particle dilemma only occurs for cases where the wavelength transforms like a propagation distance. Material waves (e.g. water or sound) do not carry this Doppler factor on $\lambda$, merely experiencing length contraction during Lorentz boosting.
We now have a resolution of our “paradox”. The de Broglie wavelength appears non-covariant if we wish it to be an actual length of a structure that exists simultaneously in space (like a wave). However, if we regard it as a propagation distance instead, with no inherent simultaneity in any frame, there is perfect accord with relativistic covariance. Thus, in using $\lambda$ as a length-scale to explain phenomena like interference and diffraction of quantum objects we are tacitly violating relativistic covariance. It is therefore not possible to ascribe actual waviness to quantum objects in a physically consistent manner, as is similarly concluded in [8] (although our argument here is different). All of this should cast some doubt on the physical, rather than mathematical, validity of de Broglie-Bohm theories, and potentially also on the correspondence between field modes and particles used in quantum field theory. A manifestation of the former may be seen in [9], where real pilot waves fail to reproduce the double slit experiment, possibly implying only those with physically inconsistent wavelengths can do so.
@@
We note some similarity to the double-slit explanation proffered in [15], where particles traverse only a single slit and do not “self-interfere”. The difference can be understood in terms of the usual question: “how does the particle know about a slit it doesn’t traverse”. Conventionally, waves are the answer. In [15], it is non-local particle properties. However, our picture is that the slits’ quantum structure places limits on the allowed local interactions with incoming particles. Thus, a two-slit structure’s difference to a single slit requires no non-local effects, unless all phenomena described by phonons in condensed matter are also deemed non-local. Additionally, it has been argued that diffraction of delocalised matter waves, on the molecular scale, requires accounting for local interactions between the molecule and diffraction element [16]. We extend this work by arguing it is true even for point particles.
@@
This paper is structured as follows: in section 2 we explore the explanation of the Kapitza-Dirac effect and how it has an elegant description in terms of local particle interactions. Then, in section 3 we explore how this description can be extended to other scenarios where diffraction and interference occur. To prove that quantised particle interactions are a viable alternative to interference we explore a double-slit experiment with photon orbital angular momentum in section 5. Then we demonstrate that quantum mechanics requires the far-field double slit pattern is determined at the slits in section 6 and discuss coherence as well as “which way” information in section 7. Finally, we draw conclusions in section 8.
@@ 2 The Kapitza-Dirac effect
This phenomenon, proposed in 1933 [17] and later confirmed [18, 19], is where an electromagnetic standing wave acts as a diffraction grating for charged (or polarisable) particles. Ostensibly this represents a symmetry between light and matter, which is a reflection of the particle-wave duality. However, the KDE has a wave-particle object diffracted by a wave, whereas conventional diffraction has a wave diffracted by a classical barrier. This seems decidedly non-dual. The standard explanation is we are discussing a phase grating, i.e. spatially varying phase differences are applied to the matter wave by the standing light wave due to the electromagnetic coupling of the particles [12]. We can immediately see that this is not actually explained by waves, both wave and particle properties are needed.
However, the presence of diffraction maxima in the KDE can be fully explained in terms of incident particles absorbing discrete light quanta of fixed energy [12]. In particular, the particle (travelling initially in the $x$-direction) absorbs a photon with $y$-direction momentum $p_{y}=\hbar k$ ($k$ is the laser wavenumber) and subsequently emits a photon at $p_{y}=-\hbar k$ via stimulated emission from the reflected laser beam. Thus, the incident particle incurs a momentum change of $2\hbar k$. Since this can happen multiple times as the particle crosses the interaction region, we would expect a momentum spectrum for the out-going particles to be given in terms of $2n\hbar k$, where $n$ is an integer. The intensity peaks are indeed found to correspond to these values.
@@ 3 Diffraction by material gratings
We turn then to more conventional diffraction phenomena, electromagnetic or matter waves diffracted by material gratings. These scenarios apparently demand a wave-like description, even if this requires that a particle’s wave aspect “self-interferes” and suffers from our subtle inconsistency with all forms of relativity. To facilitate our search for a picture where the diffraction element is no longer a passive, classical barrier, we will consider a 1D diffraction grating, centred at $x=0$ and $y=0$, that extends symmetrically in the $y$ direction. A distant screen is used to detect diffracted particles at $x=L$, where $L\gg d$ (the grating periodicity scale). It is well known that the diffraction maxima of a periodic grating can be found at angles $\theta$ (measured from the $x$-axis) given by
@@
where $m$ is an odd integer. Thus, the forbidden scattering angles, that apparently required a wave-based explanation, can be understood as being ruled out by requiring a forbidden half-quantum exchange of momentum. To elaborate, the $y$-momentum needed for deflection of the particle to “destructive interference” positions cannot be obtained from interactions with the periodic scatterer or its constituents, due to quantisation of collective momentum states (to be justified below). This quantisation rule was first observed by Duane [11] in 1923, but has only received scattered mentions in the literature over the last century [20, 13, 21, 6, 4, 22, 23, 24]. The generalised version of the simple analysis above fully reproduces both Laue and Ewald conditions [21]. Notably, this viewpoint requires that the diffraction grating must be characterised as inherently quantum. However, this comes with a difficulty: how to justify the quantised nature of the interaction? Mathematical justification is easy to find [22, 23]: Bloch’s theorem [25] requires that structures characterised by spatially periodic potentials have quantised momentum eigenstates. For a crystalline solid, these are collective motion states or phonons. The scattering interaction could therefore be similar to the physics surrounding the Mössbauer effect [26], where the incident particle interacts with a local lattice site and recoil can be absorbed by collective motion [27, 28]. The interactions that mediate collective motions naturally restrict local recoils as well, hence the points of “destructive interference” corresponding to forbidden lattice momenta. Thus, the intensity peaks correspond to resonant scattering, where exact lattice quanta are exchanged, and the continuity is supplied by local recoils. Notably, the explanation of the Mössbauer effect eliminates the problem that stymied Landé [6]: it is possible for a local interaction to depend upon the global structure of a participant. Importantly, this explanation hinges upon the periodic structure of the diffraction grating, with Bloch’s theorem assuring us it will extend to any periodic diffraction phenomenon and [22] extending it to less symmetric structures. However, there is no reason that the same physical mechanism should explain every example of such interference. Symmetry guarantees interference mathematically, but does not uniquely specify any physics.
@@
The looming question is whether such a model can be compatible with the uncertainty principle, as it is commonly held that a particle cannot possess definite momentum and position at the same time (see e.g. [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 1]). This is not a problem for Duane’s model, as the aforementioned belief is a fallacious use of the principle [4, 2] with many more conventional texts acknowledging this [40, 41, 42, 43, 44]. As rigorously defined, the uncertainty principle applies to standard deviations over a measurement ensemble [45], it thus has nothing to say about the ontic status of individual particle properties. Indeed, Popper argued that were the principle to apply to individual particles, the minimum ensemble deviation could never be probed [2].
@@ 4 Evidence supporting Duane’s model
Interestingly, there are experiments that already exhibit a hallmark of Duane’s mechanism: the sensitivity of diffraction patterns to excitation of phonon modes within the diffraction element prior to scattering [46, 47, 48]. The explanation of these results is explicitly done in terms of particle-phonon scattering interactions but is often written in “wave language”. For instance, in [46, 47], the authors describe the exchanged wave number $Q$ by $Q=q+K_{Q}$, where $q$ is the phonon wave number and $K_{Q}$ is the closest reciprocal lattice vector to $Q$. Thus, when no phonons are excited, the exchange is quantised according to the reciprocal lattice vector, which follows from the general form of Eq. (4) [21, 22]. When photons are excited, the situation is very closely analogous to the KDE, we have an incident particle diffracted by a phonon wave, and we cannot explain this by resorting to wave-wave diffraction without referring to particle interactions. However, we have the option of a complete, physically consistent, description in terms of local particle-particle interactions. Here the phonon is understood as a representation of quantised states of collective motion enforced by the lattice symmetries, interaction with such states must then be mediated via local interactions at lattice sites. Thus, advanced diffraction experiments with practical applications already use Duane’s picture without even realising it.
@@
A further point of interest arises in [47]. Here, it is demonstrated that exciting phonons alters the diffraction spectrum in a very particular manner. The apparent effect is that phonon excitation broadens the Bragg peaks normally associated with the target crystal. The more rapidly the probe follows the excitation, the more pronounced this effect is. What this demonstrates is a case where continuity is supplied between otherwise discrete peaks. This is important, as it demonstrates an analogy to precisely the question we should ask about Duane’s mechanism: why is the spectrum continuous if the momentum states are strictly quantised? While it is unlikely to be the same mechanism that answers the question in general, it does suggest that physical processes exist that perform the required role.

### v4

@@ Abstract
The dual wave-particle nature of quantum objects is a notoriously unintuitive feature of quantum theories. However, it is often deemed essential, due to quantum objects exhibiting diffraction and interference. We extend the work of Landé and Lévy-Leblond to demonstrate that de Broglie wavelengths are not relativistically covariant as simultaneous spatial structures, making wave properties an unviable explanation of apparent interference. We then explore whether modern experiments vindicate an alternative view: that apparent waviness in diffraction and interference scenarios emerges as a consequence of quantised interactions between particles. Such a view has historically received very little attention, despite being the exact modern explanation of both the Kapitza-Dirac effect and ultrafast electron diffraction. We then study a photon orbital angular momentum realisation of the double slit to show that quantised exchanges can mimic interference. Finally, we demonstrate that the quantum formalism demands that a consistent momentum spectrum, determined at the point of scattering, undergirds double-slit interference patterns. This cannot be reconciled with the idea that particle momentum is determined by wave overlap beyond the slits.
@@ 1 Introduction
There is, however, a far graver problem with quantum waviness: the wavelengths are not covariant as correlation lengths in any form of relativity. In 1975 Landé argued [6] that because de Broglie wavelengths take the form $\lambda=\frac{h}{p}$ they must vary with inertial frame. Landé also assumed that all real waves should possess phase invariance, generating an apparent paradox. Lévy-leblond [7] showed that $\lambda$ varying with inertial frame is necessary to maintain the Galilean invariance of Schrödinger equation solutions and thus phase invariance does not hold for quantum waves. This establishes that $\lambda=\frac{h}{p}$ is not a Galilean invariant, but sheds no light on why this is the case. We demonstrate that $\lambda$ scales like the distance propagated by a particle, for both Galilean and Lorentzian cases. By this we mean that we cannot measure $\lambda$ as the simultaneous length of an extended structure in space. It thus cannot be invoked as a length scale to explain interference phenomena.
@@
1.1 Contribution to the literature
Our work here provides novel arguments against the physical reality of quantum waviness. Both in terms of explaining the physics behind the de Broglie wavelength’s lack of Galilean invariance, and in a derivation of the double slit momentum distribution at the point of scattering. We also bring together experiments with photon orbital angular momentum, ultrafast electron diffraction, coherence, momentum conservation, and the impact of polarisability on interference patterns to build an original case for Duane’s model of quantum diffraction and interference. Additionally, a novel analysis of the quantum eraser highlights the power of this viewpoint in resolving mysterious phenomena and explaining why wave explanations arise in the first place. We believe this constitutes a new and unanswerable case against quantum interpretations that invoke physical wave properties, while simultaneously arguing for the reconsideration of a side-lined hypothesis.
1.2 Paper structure
@@
This paper is structured as follows: in section 2 we demonstrate that de Broglie wavelengths are not associated with a simultaneously measurable length scale. Then, in section 3 we explore the description of the Kapitza-Dirac effect in terms of local particle interactions. In section 4 we explore how this particle description can be extended to other scenarios where diffraction and interference occur. To prove that quantised particle interactions are a viable alternative to interference, we explore experiments that support the position in section 5. We next demonstrate that quantum mechanics requires the far-field double slit pattern is determined prior to wave overlap in section 6 and discuss “which-way” information and quantum erasers in section 7. Finally, we draw conclusions in section 8.
2 The problem of wavelengths
As has been mentioned already, de Broglie wavelengths are not Galilean invariants. This is usually dismissed, as there exist covariant derivations of the de Broglie relation. However, it should cause us pause, a wavelength requires an extended structure, defined simultaneously at multiple points in space. If a de Broglie $\lambda$ is not a Galilean invariant, can it actually be such a length?
Relativistic quantum theories use $p^{\mu}=\hbar k^{\mu}$ as a definition of a covariant wave-vector. This can be shown to lead directly to the de Broglie relation. Thus, we can provide a covariant mathematical description, but the wavelength is not a Galilean invariant. This can only mean that it does not merely scale with $1/\gamma$ under Lorentz boosting. The dependence on relative velocity must be different. This difference should allow us to determine what $\lambda$ actually is.
Thus, we consider the Lorentz momentum transform for a particle moving at $u$ in the same direction as the boost $v$
$p^{\prime}=\gamma\left(p-v\frac{E}{c^{2}}\right)=\gamma p\left(1-\frac{v}{u}\right)\;,$ (1)
the de Broglie wavelength is then
$\lambda^{\prime}=\frac{h}{\gamma p\left(1-\frac{v}{u}\right)}=\frac{\lambda}{\gamma}\frac{u}{u-v}\;.$ (2)
This is consistent with the Galilean limit when $\gamma\to 1$, the usual length transform is modified by the factor $\frac{u}{u-v}$.
Now, the Lorentz transform for distances propagated by an object with speed $u$
$\Delta x^{\prime}=\gamma\Delta x\frac{u-v}{u}\;.$ (3)
The structure is remarkably similar to the transform for the de Broglie wavelength. Thus, the lack of Galilean invariance arises because $\lambda$ is not an actual length of a structure that exists simultaneously in space (like a wave). The relativistic covariance matches that of a propagation distance instead, with no inherent simultaneity in any frame. Thus, we cannot interpret $\lambda$ as a length-scale which could explain phenomena like interference and diffraction without tacitly violating relativistic covariance. Furthermore, as there is no simultaneity to de Broglie waves, they are not waves. Actual wave phenomena do not suffer from this, their wavelengths are lengths and can be measured simultaneously.
There is a snag, however. Electromagnetic waves obey the de Broglie relation and yet evidently have spatial extent, simultaneity, and measurable amplitudes. Does this demonstrate a failure of the argument above? No, as such a wave is not one, but a stupendous number of photons. Each photon is not a wave, but the above argument does not necessarily apply to their collective. Each may be thought of as a soliton-like pulse. A cascade of such pulses, correlated by their source, gives us what we need to build a real wave. Such wave-like coherence can indeed appear in the classical limit [13] and be analysed in terms of interactions, rather than “waviness” [14].
It is therefore not possible to ascribe actual waviness to quantum objects in a physically consistent manner, as is similarly concluded in [15] (although our argument here is novel). All of this should cast some doubt on the physical, rather than mathematical, validity of de Broglie-Bohm theories, and potentially also on the correspondence between field modes and particles used in quantum field theory. A manifestation of the former may be seen in [16], where real pilot waves fail to reproduce the double slit experiment, possibly implying only those with physically inconsistent wavelengths can do so.
@@ 3 The Kapitza-Dirac effect
This phenomenon, proposed in 1933 [17] and later confirmed [18, 19], is where an electromagnetic standing wave acts as a diffraction grating for charged (or polarisable) particles. Ostensibly this represents a symmetry between light and matter, which is a reflection of the particle-wave duality. However, the KDE has a wave-particle object diffracted by a wave, whereas conventional diffraction has a wave diffracted by a classical barrier. This seems decidedly non-dual. The standard explanation is we are discussing a phase grating, i.e. spatially varying phase differences are applied to the matter wave by the standing light wave due to the electromagnetic coupling of the particles [10]. We can immediately see that this is not actually explained by waves, both wave and particle properties are needed (unlike conventional optical diffraction).
The presence of diffraction maxima in the KDE can be fully explained in terms of incident particles absorbing discrete light quanta of fixed energy [10]. In particular, the particle (travelling initially in the $x$-direction) absorbs a photon with $y$-direction momentum $p_{y}=\hbar k$ ($k$ is the laser wavenumber) and subsequently emits a photon at $p_{y}=-\hbar k$ via stimulated emission from the reflected laser beam. Thus, the incident particle incurs a momentum change of $2\hbar k$. Since this can happen multiple times as the particle crosses the interaction region, we would expect a momentum spectrum for the out-going particles to be given in terms of $2n\hbar k$, where $n$ is an integer. The intensity peaks are indeed found to correspond to these values.
@@ 4 Diffraction by material gratings
We turn then to more prosaic diffraction phenomena: electromagnetic or matter waves diffracted by material gratings. These scenarios apparently demand a wave-like description, even if this requires that a particle’s wave aspect suffers from our subtle inconsistency with all forms of relativity.
4.1 Quantisation of momentum
To facilitate our search for a picture where the diffraction element is no longer a passive, classical barrier, we will consider a 1D diffraction grating, centred at $x=0$ and $y=0$, that extends symmetrically in the $y$ direction. A distant screen is used to detect diffracted particles at $x=L$, where $L\gg d$ (the grating periodicity scale). It is well known that the far-field diffraction maxima of a periodic grating can be found at angles $\theta$ (measured from the $x$-axis) given by
@@ 4.1 Quantisation of momentum
Thus, the forbidden scattering angles, that apparently required a wave-based explanation, can be understood as being ruled out by requiring a forbidden half-quantum exchange of momentum. To elaborate, the $y$-momentum needed for deflection of the particle to “destructive interference” positions cannot be obtained from interactions with the periodic scatterer or its constituents, due to quantisation of collective momentum states (to be justified below). This quantisation rule was first observed by Duane [9] in 1923, but has only received scattered mentions in the literature over the last century [20, 11, 21, 6, 4, 22, 23, 24]. The generalised version of the simple analysis above fully reproduces both Laue and Ewald conditions [21]. Notably, this viewpoint requires that the diffraction grating must be characterised as inherently quantum, removing the artificial separation of quantum and classical within the experiment.
4.2 Bloch’s theorem and the Mössbauer effect
Duane’s model does come with a difficulty: how to justify the quantised nature of the interaction? Mathematical justification is easy to find [22, 23] in Bloch’s theorem [25], which requires that structures characterised by spatially periodic potentials have quantised momentum eigenstates. For a crystalline solid, these are particular collective motion states or phonons. The scattering interaction could therefore be similar to the physics surrounding the Mössbauer effect [26], where the incident particle interacts with a local lattice site and recoil can be absorbed by collective motion [27, 28]. The interactions that mediate collective motions naturally restrict local recoils as well, hence the points of “destructive interference” corresponding to forbidden lattice momenta. Thus, the intensity peaks correspond to resonant scattering, where exact lattice quanta are exchanged, and the continuity is supplied by local recoils or a low energy background of excited phonon states.
Notably, the explanation of the Mössbauer effect eliminates the problem that stymied Landé [6]. It is possible for a local interaction to depend upon the global structure of a participant. Importantly, this explanation hinges upon the periodic structure of the diffraction grating, with Bloch’s theorem assuring us it will extend to any periodic diffraction phenomenon and [22] extending it to less symmetric structures. However, there is no reason that the same physical mechanism should explain every example of such interference. Symmetry guarantees interference mathematically, but does not uniquely specify any mediating physics.
4.3 Does this conflict with the uncertainty principle?
@@
The looming question is whether such a model can be compatible with the uncertainty principle, as it is commonly held that a particle cannot possess definite momentum and position at the same time (see e.g. [29, 30, 31]). This is not a problem for Duane’s model, as the aforementioned belief is a fallacious use of the principle [4, 2] with several popular texts acknowledging this [32, 33, 34, 35]. As rigorously defined, the uncertainty principle applies to standard deviations over a measurement ensemble [36], it thus has nothing to say about the ontic status of individual particle properties. Indeed, Popper argued that were the principle to apply to individual particles, the minimum ensemble deviation could never be probed [2].
@@ 5 Evidence supporting Duane’s model
Interestingly, there are experiments that already exhibit a hallmark of Duane’s mechanism as well as effects that undermine wave explanations. These will be detailed below.
5.1 Bragg diffraction
When X-rays are incident on a crystal lattice they are well-known to experience preferential reflection to discrete angles. Interestingly, electrons at the same energies experience a somewhat different behaviour. Allegedly, both of these particles correspond to the waves of the same wavelength. However, despite experiencing the same boundary conditions, these waves experience different diffraction/interference behaviours [37]. Particularly, the electrons produce many more diffraction spots than X-rays, and these spots are approximately patterned after the reciprocal lattice of the scatterer. The explanation is simple: electrons interact far more strongly with the crystal than photons do. However, this already betrays the wave explanation. It is the particle interactions that determine the resultant pattern, and the quantisation of diffraction angles is determined by the reciprocal lattice vector. This is precisely what is implied by the more general form of Eq. (5), as demonstrated in [21] which also derives the Laue and Ewald conditions. In addition, the reciprocal lattice projection is the momentum representation of the crystal’s periodicity and only occurs perpendicular to the electron beam direction for transmission-type experiments. The electrons are thus gaining quantised momenta, determined in both magnitude and direction by the lattice periodicity.
5.2 Ultrafast electron diffraction
One expectation of Duane’s model would be sensitivity of diffraction patterns to excitation of phonon modes within the diffraction element prior to scattering. The presence of excited phonon modes, that might be negligible in ordinary conditions for the target, will change the possible interactions with the scatterer. This being consistent with the assertion that all diffraction is a consequence of interactions with the incident particles. One might be tempted to invoke a wave and particle explanation here, waves result from boundary conditions and the phonon interactions contribute a distinct extra component. However, this is untenable given existing experiments [38, 39, 40].
In [39] it is demonstrated that exciting phonons alters the diffraction spectrum in a very particular manner. The apparent effect is that phonon excitation shifts intensity away from the Bragg peaks to a surrounding continuum. The more rapidly the probe follows the excitation, the more pronounced this effect is. Thus, there are not two separate effects: wave interference and phonon interactions. In fact, the phonon interaction appears to suppress the distinct interference peaks [39]. This makes sense if we are discussing competing interactions, something strongly suggested by the form of the wavenumber being exchanged between particle and scatterer $Q=q+K_{Q}$ [38], where $q$ is the phonon wavenumber and $K_{Q}$ is the closest reciprocal lattice vector to $Q$. This is very clearly a momentum quantisation condition. When no phonons are excited, the exchange is quantised according to the reciprocal lattice vector, which follows from the general form of Eq. (5[21, 22]. When a significant phonon background is present, they supply the additional momentum $q$ to disperse the strict Bragg quantisation. Notably, the final pattern depends upon both the phonon energies and their relative population. These phonons need not be viewed as waves, as they represent collective lattice motions realised by local sites.
@@ 5.2 Ultrafast electron diffraction
Additionally, the phonon spectrum prior to excitation also plays a role [39]. Given that phonon modes distributed intensity away from the quantised Bragg peaks, this demonstrates is a case where continuity is supplied between otherwise discrete peaks, even in an unexcited target. We thus have a possible mechanism answering precisely the question we should ask about Duane’s mechanism: why is the spectrum often continuous if the momentum states are strictly quantised? While it is unlikely to be the same mechanism that answers the question in general, it does suggest that physical processes exist that perform the required role.
5.3 Diffraction of polarisable molecules